If it has rained 276 mm in the last 3 days, how many cm of rain will fall in one full week (7 days)? Assume that the rain continues to fall at the same rate.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Measure of rain in 3 days = 276 mm
Measure of rain in one day = 276 / 3
= 92 mm
So, measure of rain in one week i.e 7 days = 92 × 7
=644 mm
= 644 / 10 = 64.4 cm

Thanks